# Fast Revenue Maximization

### v1

@@ 3 Applications to Price Experimentation
We have established a tractable procedure to evaluate the value of any historical information set that comes in the form of prices and their associated conversion rates. In the following subsections, we illustrate how Theorem 1 can be leveraged to derive insights on price experimentation. In Section 3.1, we study the value of local experimentation and consider a setting in which the decision-maker cannot substantially vary prices and therefore has only an estimate of the gradient at a given price along with the conversion rate at that price. In Section 3.2, we investigate the value of global experimentation and derive an optimal (freely selected) second price to experiment at after receiving the conversion rate from a first price. Finally, in Section 3.3, we illustrate how our exact analysis could be used to considerably reduce the number of price experiments needed to reach a certain level of optimality, when using classical algorithms.
@@ 3.1 Pricing with gradient information: the value of local experimentation
In many practical settings, the seller cannot drastically change the offered price and therefore only gather local informations. Motivated by this, we consider the setting in which the seller knows the support of the demand function $[\underline{v},\overline{v}]$ and has observed two demand points, i.e. $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{1},q_{1}),(p_{\epsilon},q_{\epsilon})\}$. The two prices are such that $p_{\epsilon}=(1+\epsilon)p_{1}$ for $\epsilon>0$. In what follows, we will parametrize this instances by the estimated gradient of the revenue function at $p_{1}$ defined as,
$\nabla_{p}Rev(p|\overline{F})(p_{1})\approx\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}:=g_{\epsilon}.$
Value of the numerical gradient. We propose to understand the value of such observations by leveraging the reduction in Theorem 1 to numerically compute the maximin ratio defined in (Maximin) when the demand function is regular (i.e., $\pazocal{C}=\pazocal{F}$) and its support is included in $[\underline{v},\overline{v}]=[1,100]$. We consider the setting where the seller experiments at an increment of $\epsilon=1\%$ of the observed price $p_{1}$ which we fix at $10$ and we consider different initial observed demand $q_{1}$ in $\{0.01,0.1,0.25,0.5,0.75,0.9,0.99\}$. For each initial $q_{1}$, we evaluate a lower bound on the maximin Ratio for a grid of 200 possible values of $g_{\epsilon}$ and report the results in Figure 2. In Appendix A, we discuss in details how we compute the lower bound on (Maximin).
@@
Figure 2: The value of a gradient measurement for various initial settings $q_{1}$.
Figure 2 naturally demonstrates that when the absolute value of the gradient is low, the worst-case ratio is guaranteed to be high as the decision-maker is very close to the optimal price. Interestingly, relatively large values of the gradient are still very informative and allow considerable improvement compared to a single price observation. When $q_{1}=0.5$, i.e., when the first price was at the median of the value distribution, the worst-case ratio without gradient information is around $60\%$ (as we will see in Figure 3). This can increase to more that $80\%$ with a gradient with an absolute value lower or equal than $0.1$.
Value of the gradient-sign: In settings where the estimation of the absolute value of the gradient is noisy and unreliable and where the seller can only estimate the sign of the gradient, our framework allows characterizing the value of such measurement when the demand is regular as it translates to the elimination of one side of the interval of feasible optimal reserve prices. Indeed, given that the revenue function is unimodal in that case, the sign of the gradient at a price $p_{1}$ allows to assess whether the optimal price ${r^{*}}$ satisfies ${r^{*}}\geq p_{1}$ or ${r^{*}}\leq p_{1}$.
We illustrate numerically the value of a positive and negative gradient measurement in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$, the seller estimates the gradient sign at the observed price $p_{1}$ which we fix at $10$, and we consider different initial observed demand $q_{1}$. For each initial setting $q_{1}$, we evaluate the maximin Ratio for negative and positive observed gradient-sign as well as the value of the maximin Ratio without gradient information and report the values in Figure 3.
Figure 3: The value of the sign of gradient measurement for various initial settings $q_{1}$.
@@
3.2 The value of one additional experiment: the value of global experimentation
We now relax the constraint that the decision-maker should explore locally and consider the setting in which the seller has observed a single demand point $(p_{1},q_{1})$ and has the opportunity to experiment at any second price $p_{2}$ by collecting the demand level at the latter price, before committing to a final pricing mechanism. We model the problem as a game between Nature and the seller, in which the seller commits to a price $p_{2}$ to experiment at, and Nature counters the seller by choosing the associated feasible quantile $q_{2}$ to yield the worst maximin ratio in the final stage. More specifically, for a fixed $p_{2}$, Nature is solving the following problem,
@@ 3.2 The value of one additional experiment: the value of global experimentation
We illustrate numerically the value of one additional experiment in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. The seller starts with one initial observed demand at the price $p_{1}=10$. For each initial setting $q_{1}$ and candidate experiment price $p_{2}$ from a fixed grid of $[\underline{v},\overline{v}]$ of size 200, we numerically evaluate a lower-bound on the above objective using the procedure described in Appendix A. We present in Figure 4-a, the improvement in the worst-case ratio allowed with the best experiment price $p_{2}$ chosen by the decision-maker and show in Figure 4-b what this best price is.
@@
b) ”Best” candidate experiment price $p_{2}$.
Figure 4: The value of one additional global experiment for various initial settings $(p_{1},q_{1})$.
Figure 4 shows that by allowing the decision-maker to select two prices (as opposed to one), one can considerably improve the performance of the pricing policy. Indeed, while a single price already guarantees a ratio of at most $30\%$ for conversion rates between $[0.01,0.99]$, the decision-maker is ensured to obtain almost $50\%$ of the revenue by experimenting correctly at a second price. Figure 4-b also illustrates how our procedure allows us to select a second price to run efficient experimentations. For instance, it provides general guidelines regarding when the price should be deflated or inflated to gather more information: in this case, when the conversion rate at $p_{1}$ is lower (respectively higher) than approximately $0.3$ the decision-maker should consider experiment more by deflating (resp. inflating) the price.
3.3 Dynamic pricing with deterministic feedback
Section 3.2 provides an initial step towards dynamic pricing, where the decision-maker explores at one more price before committing. In this section, we use our framework to provide a criterion to stop pricing experiments while guaranteeing a certain level of performance. This problem is of significant importance as price experimentation is particularly costly for firms: ending them as soon as possible can lead to substantial savings. To illustrate this, we consider the case where the seller is dynamically adjusting the posted prices and observes deterministic true demand values at the implemented prices in each round.
Consider a general dynamic pricing algorithm $\theta$ that maps the history $\pazocal{H}_{t}$ of collected prices and associated observed deterministic demands in each iteration $t$ to the next price to post i.e $p_{t}=\theta\left(\pazocal{H}_{t}\right)$. We propose to use our framework as a stopping criterion for the price experimentation phase. More specifically, we propose to augment any dynamic pricing method $\theta$ as in Algorithm 1.
@@ 3.3 Dynamic pricing with deterministic feedback
Initialize $\pazocal{H}_{1}$, t =1;
while $\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\pazocal{H}_{t})\right)>\epsilon$ do
       Offer price $p_{t}=\theta(\pazocal{H}_{t})$ posted by algorithm $\theta$ ;
       Update $\pazocal{H}_{t+1}=\pazocal{H}_{t}\cup\{(p_{t},q_{t}=\overline{F}(p_{t}))\}$ and increment $t$;
@@
Return $p_{t^{*}},$ where $t^{*}=\arg\max_{t}\left(p_{t}\cdot q_{t}\right)$;
Algorithm 1 Revenue Maximization Algorithm $\theta$ with stopping criterion
When the underlying value distribution is regular, i.e., $\pazocal{C}=\pazocal{F}$, the revenue function is unimodal. In this case, a popular algorithm is the Ternary search algorithm (described in Appendix B). We propose to illustrate the effectiveness of our proposed framework in reducing the number of price queries needed to ensure that the worst-case ratio is at least $99\%$ and compare different stopping criteria combined with the Ternary search algorithm.
In particular, we sample random instances from a parametric family of demand curves (either linear demand cures or exponential ones) and, for each instance, report the number of price queries needed to achieve 99% Ratio performance when $i)$ using the common stopping criterion in Ternary Search, $ii)$ using our criterion with $\pazocal{C}=\pazocal{F}$, i.e. when we know that the demand is regular and $iii)$ using our criterion with $\pazocal{C}=\pazocal{G}$, i.e., without using the fact the demand is regular. We set $[\underline{v},\overline{v}]=[1,100]$ (see implementation details in Appendix B). We report in Figure 5 our results.
a) Linear underlying demand.
b) Exponential underlying demand.
@@
Figure 5: Number of price queries needed to achieve 99% Ratio performance.
@@
Figure 5 shows that using our characterization of the worst-case ratio allows to stop considerably earlier than using standard stopping rules with Ternary search. When using our criterion without including the information that the demand is regular, we are able to divide by two the number of price queries needed on average compared to the usual bounds. Furthermore, by incorporating the regularity assumption in our criterion we can reduce by a factor $4$ the query complexity. We also note that our stopping criterion is data-driven and changes as a function of the observed conversion rates: for some instances, we stop after only 5 price queries (as opposed to $44$ with usual rules).
@@ 4 Conclusion
We present in this work a framework to characterize the value of data obtained through price experiments. We use this framework both to quantify the value of diverse type of historical data but also to prescribe efficient ways of running price experiments. Throughout this work, we showed that local experiments (such as the sign or the value of the gradient at an offered price) may already contain a lot of information. We also illustrated how our framework can be used to select the next price at which one should experiment and the value associated with such decisions. Finally, our analysis allows to significantly reduce the length of the experiment even when committing to standard dynamic algorithms (such as Ternary search) by providing a data-driven stopping criterion.
We believe that this work opens up many exciting research avenues. One natural question would be to extend our result to be applicable to quantify the value of noisy observations of the demand. Furthermore, the use of our method to prescribe the last price at which to experiment before committing to a revenue maximizing price (as discussed in Section 3.2) provides an interesting first step towards more efficient algorithms trading-off exploration and exploitation in dynamic pricing.

### v2

@@ 4 Applications to Price Experimentation
We described in Section 3 a tractable procedure to quantify the value of any historical information set that comes in the form of prices and their associated conversion rates. In this section, we illustrate how our analysis can be leveraged to derive insights on price experimentation. In Section 4.1, we study the value of local experimentation and consider a setting in which the decision-maker cannot substantially vary prices and therefore has only an estimate of the gradient at a given price along with the conversion rate at that price. In Section 4.2, we investigate the value of global experimentation and derive an optimal (freely selected) second price to experiment at after receiving the conversion rate from a first price. Finally, in Section 4.3, we illustrate how our exact analysis could be used to considerably reduce the number of price experiments needed to reach a certain level of performance guarantee.
@@ 4.1 Pricing with gradient information: the value of local experimentation
In many practical settings, the seller cannot drastically change the offered price and therefore can only gather “local” information. Motivated by this, we consider the setting in which the seller knows the support of the demand function $[\underline{v},\overline{v}]$ and has observed two demand points, i.e. $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{1},q_{1}),(p_{\epsilon},q_{\epsilon})\}$. The two prices are such that $p_{\epsilon}=(1+\epsilon)p_{1}$ for $\epsilon>0$.
Value of the numerical gradient. We propose to understand the value of such observations by leveraging the procedure described in Section 3 to numerically compute the maximin ratio defined in (Maximin) when the demand function is regular (i.e., $\pazocal{C}=\pazocal{F}$) and its support is included in $[\underline{v},\overline{v}]=[1,100]$. We consider the setting in which the seller experiments at an increment of $\epsilon=1\%$ of the observed price $p_{1}$ which we fix at $10$ and we consider different initial observed demand $q_{1}$ in $\{0.01,0.1,0.25,0.5,0.75,0.9,0.99\}$. For each initial $q_{1}$, we evaluate a lower bound on the maximin ratio for a grid of 200 possible values of $q_{\epsilon}$ and report the results in Figure 2. For the sake of interpretability, the x-axis of Figure 2 is parametrized by $g_{\epsilon}$ which is defined as,
$g_{\epsilon}=\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}.$
We interpret $g_{\epsilon}$ as the gradient of the revenue curve at $p_{1}$ given that, $\nabla_{p}\mathrm{Rev}(p_{1}|F)\approx\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}$.
@@
Figure 2: Value of a gradient measurement for various $q_{1}$ ($\pazocal{C}=\pazocal{F}$). The maximin ratio is computed by evaluating $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ with a discretization involving $M=2500$ points.
Figure 2 naturally demonstrates that when the absolute value of the gradient is low, the worst-case ratio is guaranteed to be high as the decision-maker is very close to the optimal price. Interestingly, relatively large values of the gradient are still very informative and allow considerable improvement compared to a single price observation. When $q_{1}=0.5$, i.e., when the first price was at the median of the value distribution, the worst-case ratio without gradient information is around $60\%$ (as we will see in Figure 3). By knowing that the absolute value of the gradient is lower or equal to $0.1$, the worst-case ratio can increase to more that $80\%$.
Value of the gradient-sign: In settings where the estimation of the absolute value of the gradient is noisy and unreliable and where the seller can only estimate the sign of the gradient, our framework allows to characterize the value of such measurement when the demand is regular as it translates to the elimination of one side of the interval of feasible optimal reserve prices. Indeed, given that the revenue function is unimodal in that case, the sign of the gradient at a price $p_{1}$ allows to assess whether the optimal price ${r^{*}}$ satisfies ${r^{*}}\geq p_{1}$ or ${r^{*}}\leq p_{1}$.
We next illustrate numerically the value of a positive and negative gradient measurement in the setting in which the seller knows that the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. As a baseline, we first consider the maximin ratio obtained by the retailer when the information set is $\{(p_{1},q_{1})\}$, i.e., when the seller knows the conversion rate at a single price. We then evaluate the maximin ratio obtained by the seller who also knows the gradient sign at the observed price $p_{1}$. For each initial conversion rate $q_{1}$, we evaluate the maximin ratio for negative and positive observed gradient-signs as well as the value of the maximin ratio without gradient information (our baseline) and report the values in Figure 3.
Figure 3: Value of the gradient-sign for various $q_{1}$. The maximin ratio with gradient-sign information is computed by modifying the evaluation of $\underline{\pazocal{L}}(\pazocal{C}(\{(p_{1},q_{1})\}),\mathbb{A})$ where we only consider the constraints in (3.3b) for $i$ such that $a_{i}\leq p_{1}$ (resp. $a_{i}\geq p_{1}$) when the gradient-sign is negative (resp. positive). ($p_{1}=10$)
@@
This example shows that depending on the initial conversion rate $q_{1}$ the marginal value associated with knowing the sign of the gradient or the numerical value of the gradient varies considerably. Our procedure allows to quantify the value of running longer and more accurate (hence more expensive) local price experiments as a function of the data at hand.
4.2 The value of global experimentation
We now relax the constraint that the decision-maker should explore locally and consider the setting in which the seller has observed a single demand point $(p_{1},q_{1})$ and has the opportunity to experiment at any second price $p_{2}$ by collecting the demand level at the latter price, before committing to a final pricing mechanism. We model the problem as a game between Nature and the seller, in which the seller commits to a price $p_{2}$ to experiment at, and Nature counters the seller by choosing the associated feasible conversion rate $q_{2}$ to yield the worst maximin ratio in the final stage. More specifically, for a fixed $p_{2}$, Nature is solving the following problem,
@@ 4.2 The value of global experimentation
We illustrate numerically the value of one additional experiment in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. The seller starts with one initial observed demand at the price $p_{1}=10$. For each initial $q_{1}$ and candidate experiment price $p_{2}$ from a fixed grid of $[\underline{v},\overline{v}]$ of size 200, we numerically evaluate a lower-bound on the above objective using the LP described in Section 3 (with $M=2500$). We present in Figure 4-a, the improvement in the worst-case ratio allowed with the best experiment price $p_{2}$ chosen by the decision-maker and show in Figure 4-b what this best price is.
@@
b)“Best” candidate experiment price $p_{2}$.
Figure 4: Value of one additional global experiment for various $(p_{1},q_{1})$.
Figure 4 shows that by allowing the decision-maker to judiciously select a second price at which to collect information, they can considerably improve the performance of the final price selected. Indeed, while a single price already guarantees a ratio of at least $30\%$ for conversion rates between $[0.01,0.99]$, the decision-maker is ensured to obtain more than $50\%$ of the revenue by experimenting correctly at a second price. More generally, across conversion rates at the initial price, the decision-maker can increase the robust performance by approximately $20\%$ through an additional well-selected price experiment. Figure 4-b also illustrates how our procedure allows us to select a second price to run efficient experimentation. For instance, it provides general guidelines regarding when the price should be deflated or inflated to gather more information: in this case, when the conversion rate at $p_{1}$ is lower (respectively higher) than approximately $0.3$ the decision-maker should consider experimenting at a lower (resp. higher) price.
4.3 Stopping criterion for dynamic pricing with deterministic feedback
In Section 4.2 we considered the case of a decision-maker who explored at one more price before committing, we illustrated how our analysis can provide an initial step towards more efficient dynamic pricing algorithms. In this section, we show how our framework can be used to provide a criterion to stop pricing experiments as soon as possible while guaranteeing a certain level of performance. More precisely, given a desired accuracy level $\epsilon$, a valid stopping criterion ensures that whenever the pricing algorithm is stopped, the pricing mechanism which is outputted will generate at least a $1-\epsilon$ fraction of the optimal revenue. We will illustrate that our framework can be used to provide a stopping criterion on top of any dynamic pricing algorithm used to gather information. To illustrate this, we consider the hypothetical case in which the seller is dynamically adjusting the posted prices and observes deterministic true conversion rates at the offered prices in each round. For instance, we consider the setting in which the value distribution is regular, i.e., $\pazocal{C}=\pazocal{F}$. In this case, given that the revenue function is unimodal, the classical Ternary search algorithm, described in Algorithm 1, can be used for dynamic pricing to ensure one approaches optimal performance.
@@ 4.3 Stopping criterion for dynamic pricing with deterministic feedback
Initialize $a=\underline{v}$, $b=\overline{v}$, $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{1}=\{(\underline{v},1),(\overline{v},0)\}$;
while $t<\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}$ ;
  // Ternary search usual stopping criterion
 do
       /* Exploration procedure of the Ternary search algorithm */
       Price at $p_{1}=a+\frac{b-a}{3}$, observe $q_{1}$ ;
       Price at $p_{2}=a+\frac{2(b-a)}{3}$, observe $q_{2}$ ;
       Record observed conversion rates $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t+1}=\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}\cup\{(p_{1},q_{1}),(p_{2},q_{2})\}$ ;
      if $p_{1}\cdot q_{1}<p_{2}\cdot q_{2}$ then
             Set $a=p_{1}$ ;
       end if
      else
             Set $b=p_{2}$ ;
       end if
      Increment $t$
@@
Return $a$ if $a\cdot q_{a}\geq b\cdot q_{b}$ else return $b$ ;
  // Exploitation price
Algorithm 1 Ternary Search
The ternary search procedure proceeds in steps, sequentially eliminating regions of the price space by leveraging the unimodality of the revenue curve. We note that any dynamic pricing algorithm designed to approach the optimal performance has a structure similar to Algorithm 1. It is defined by a stopping criterion, an exploration procedure and a final exploitation price. In what follows, we fix the exploration procedure of a given dynamic pricing method $\theta$ and combine it with a stopping criterion and an exploitation price designed by leveraging our framework. Recall $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ defined in (3.3). Algorithm 2 formalizes our meta-algorithm.
Input: Dynamic pricing algorithm $\theta$; Acceptable sub-optimality $\epsilon$; parameter $M$ for discretization in (3.3);
Initialize $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{1}$;
while $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}),\mathbb{A})<1-\epsilon$ do
       Use the exploration procedure of $\theta$;
       Update $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t+1}$ and increment $t$;
end while
Return $(\psi^{*}_{i})_{i\in\{1,\ldots,M\}}$ solution of the discretized problem (3.3) ;
Algorithm 2 Meta-algorithm for dynamic pricing
Remark that Algorithm 2 provides a valid stropping criterion by design. Indeed, $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}),\mathbb{A})$, defined in (3.3), is a lower bound on the maximin ratio. Hence, whenever the algorithm stops, the maximin ratio is at least $1-\epsilon$ and the mechanism returned by (3.3) is ensured to achieve that ratio. We show in Lemma 8 (stated and proved in Section 6) that the usual stopping criterion of the Ternary search algorithm is also valid.
In what follows, we illustrate the effectiveness of our proposed framework in reducing the number of price queries needed to ensure that the worst-case ratio is at least $99\%$. To compare different stopping criteria, we fix the Ternary search algorithm structure, and vary the stopping criterion. In particular, we sample random instances from a parametric family of regular demand curves (either linear demand curves or exponential ones). The linear model satisfies for every $v\in[0,\infty)$,
$\overline{F}(v)=\min\left(\max(a-b\cdot v),0),1\right),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[1,b]$. For the exponential demand we have for every $v\in[0,\infty)$,
$\overline{F}(v)=\min(\exp(a-b\cdot v),1),$
For each instance, we report the number of price queries needed to achieve at worst-case ratio greater or equal to $99\%$ when
  1. (i)
    using the stopping criterion of Ternary search which uses $N^{\mathrm{Ternary}}(\epsilon)=2\cdot\left\lceil\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}\right\rceil$ price experiments, and equal to $44$ in this setting,
  2. (ii)
    using our criterion with $\pazocal{C}=\pazocal{F}$, i.e., when we know that the demand is regular,
  3. (iii)
    using our criterion with $\pazocal{C}=\pazocal{G}$ and adding the constraint that the revenue function is unimodal (this can be implemented by only including constraints in (3.3b) for $i\in\{1,\ldots,M+1\}$ such that $a_{i}\in[a,b]$, where $a$ and $b$ are the variables used by Algorithm 1).
We report in Figure 5 the number of price queries needed by each criterion to ensure that the exploitation price will achieve a $99\%$ worst-case ratio. We emphasize that while the stopping criterion changes, the dynamic pricing algorithm used for exploration is always the same and corresponds to the Ternary search approach.
a) Linear demand.
b) Exponential demand.
@@
Figure 5: Distribution of the number of price queries needed to achieve 99% worst-case ratio performance. ($[\underline{v},\overline{v}]=[1,100]$)
@@
Figure 5 shows that using our characterization of the worst-case ratio allows to stop considerably earlier than by using the standard stopping criterion of Ternary search. When using our stopping criterion which only uses the unimodality of the revenue curve, we are already able to divide by two the number of price queries needed on average. This difference can be explained by the data-driven nature of our criterion. Indeed, the Ternary search stopping criterion is non-adaptive and considers the worst-case number of iterations needed across all unimodal curves, whereas our stopping criterion leverages the information gathered at each step to rule out some of the unimodal curves based on the conversion rates observed.
Furthermore, our stopping criterion can be used to incorporate additional beliefs about the demand distribution. For instance, we can incorporate regularity by setting $\pazocal{C}=\pazocal{F}$. In that case, we see that our criterion can reduce on average the number of price points at which we experiment by a factor of five compared to the Ternary search criterion. For some of the instances considered our regular stopping criterion proposes to stop after only 5 price queries (as opposed to $44$ with usual stopping rules and ternary search). This significant gain in efficiency considerably changes our understanding of the value of dynamic price experiments.
@@ 5 Conclusion
We present in this work a framework to precisely quantify the value of data obtained through price experiments in the form of past offered prices and observed conversion rates. Our main methodological contribution is a reduction of an optimization problem whose value quantifies the worst-case revenue gap incurred by a decision-maker who only has access to certain conversion rates compared to an oracle knowing the full value distribution. We leverage this reduction to derive insights on the value of diverse types of historical data but also to prescribe efficient ways of running price experiments.
Throughout this work, we showed that local experiments (such as the sign or the value of the gradient of the revenue function at an offered price) may already encode information useful for pricing purposes. We also demonstrate that our exact quantification of the value of these different types of information provides a tool for rigorous cost-benefit analyses. Indeed, decision-makers may assess whether the benefits associated with obtaining a precise value of the gradient compared to its sign justify the cost of running a longer experiment.
We also illustrate how our framework can be used to improve price experiments. We first propose a procedure to select the next price at which one should experiment before committing to a final price and we quantify the value associated with such decisions. Finally, we show that our framework allows to significantly reduce the length of price experiments even when committing to standard dynamic algorithms (such as Ternary search) by providing a data-driven stopping criterion tailored to the structure of the pricing problem. This stopping criterion is adaptive to the the information collected.
We believe that this work opens up many exciting research avenues towards more efficient price experimentation. One natural question would be to extend our result to be applicable to quantify the value of noisy observations of the demand. Furthermore, the use of our method to prescribe the last price at which to experiment before committing to a final price (as discussed in Section 4.2) provides an interesting first step towards more efficient algorithms for dynamic pricing. We hope that our work inspires new algorithmic principles to trade-off exploration and exploitation in dynamic pricing over multiple steps.
Acknowledgment
The authors would like to express their sincere gratitude to Amine Allouah for his invaluable feedback and insightful discussions during the early stages of this work.